OpenStudy (anonymous):

Solve for x \[\frac{ x }{ (x+1)(x-2) }=1\]

OpenStudy (lgbasallote):

simplify your denominator first \[\implies \frac{x}{x^2 + x - 2x - 2} = 1\] simplify... \[\implies \frac{x}{x^2 - x - 2} = 1\] right?

OpenStudy (anonymous):

ok, I thought it was already simplified?

OpenStudy (lgbasallote):

by simplify i meant expand

OpenStudy (anonymous):

right. please continue.

OpenStudy (lgbasallote):

now cross multiply

Parth (parthkohli):

You could also have started with\[x = (x + 1)(x - 2)\]\[ x = x^2 - x - 2\]\[ x^ 2 - 2x - 2 = 0\]This is a quadratic equation.

OpenStudy (lgbasallote):

what do you get if you cross multiply \[\frac{x}{x^2 - x - 2} = 1\] ?

OpenStudy (anonymous):

oh I get it.

OpenStudy (anonymous):

cheers guys.

Parth (parthkohli):

@Hollywood_chrissy Lord LG tells you to multiply \(x^2 - x - 2\) to both sides.

Parth (parthkohli):

G'job

OpenStudy (lgbasallote):

if G'job means good job... does that mean B'job means bad job?

Parth (parthkohli):

...

Parth (parthkohli):

@lgbasallote you're very punny.

OpenStudy (lgbasallote):

you'll understand when you're older

Parth (parthkohli):

I know what b'job is.

OpenStudy (zzr0ck3r):

lol